OpenStudy (anonymous):

If a Circle with a 10 inch diameter has 10 equal slices with a interior angle of 60*, what would the arc length be for one slice?

OpenStudy (anonymous):

 Sorry Imeant 8 equal slices

OpenStudy (shane_b):

The diameter of the circle is 10in so the radius is 5in. First, convert 60 degrees to radians:\[(60)(\frac{\pi}{180})=1.047\]This value is then used in this equation:\[\theta_{radians}=\frac{s}{r}\]where s is the arc length and r is the radius. Therefore:\[1.047_{radians}=\frac{s}{5}\]Solving for s you get: \[s=5.235in\]

OpenStudy (cwrw238):

or you could precede as follows circumference = 10pi - this is split up into 8 equal lengths so length of arc of each slice = 10 pi / 8

OpenStudy (shane_b):

<slaps forehead> Yep.

OpenStudy (shane_b):

I think I jumped right into the formula and overt-hought it :P

OpenStudy (shane_b):

hmm...but it didn't come out the same. I must be in error somewhere up there.

OpenStudy (cwrw238):

yea - i've done something like that in past as well

OpenStudy (cwrw238):

10pi/8 = 3.927

OpenStudy (shane_b):

10pi/8 is obviously right...just not sure why the formula didn't come out the same.

OpenStudy (cwrw238):

why 60(pi/180)?

OpenStudy (shane_b):

It has to be in radians

OpenStudy (cwrw238):

yes but why 60?

OpenStudy (shane_b):

The interior angle of 60 degrees

OpenStudy (cwrw238):

shouldn't that be 45?

OpenStudy (shane_b):

I see the problem....the question states 60 degrees but it should have been 45. Yep.

OpenStudy (shane_b):

Now all is right in the world again.

OpenStudy (cwrw238):

yes - I hadn't noticed that )in problem) - confusing

OpenStudy (cwrw238):

yes - lol

OpenStudy (shane_b):

I guess I paid too much attention to the wording and not the picture :)

OpenStudy (shane_b):

And NOW I realize that it should be 36 degrees....IF you go by the wording of TEN equal slices

OpenStudy (cwrw238):

yes - and i did the opposite - questioner seems to have lost interest - he's gone!!

OpenStudy (cwrw238):

right!!!!

OpenStudy (shane_b):

That question is all jacked up :)

OpenStudy (cwrw238):

a mess!!